OpenStudy (anonymous):

A bank account pays an interest of 4.4% compounded continuously. How long does it take the amount of money in the account to triple?

OpenStudy (anonymous):

lol no bank compounds continuously, but no matter your job is to solve \[e^{0.044t}=3\] for \(t\) do you know how to do that?

OpenStudy (anonymous):

yes,

OpenStudy (anonymous):

k then that is all use a calculator to get the numeric answer

OpenStudy (anonymous):

In(3)=0.44t In(3)/0.44t t=In(3)/0.44 t=25 answer: 25years??? Is that right?

OpenStudy (anonymous):

your decimal is in the wrong place i think \(4.4\%=.044\)

OpenStudy (anonymous):

oh but 25 is right!

OpenStudy (anonymous):

you must have computed it correctly, just wrote it incorrectly you have it , 25 years

OpenStudy (anonymous):

opps! yeah i typed that wrong

OpenStudy (anonymous):

k, you are done good job hope it is clear why you set it equal to 3

OpenStudy (anonymous):

thank you!! :D

OpenStudy (anonymous):

yw

OpenStudy (tkhunny):

@satellite73 I want to start a new bank, just so I can credit continuous interest. Are you in?

OpenStudy (anonymous):

only if it is \(100\%\) a year so i can finally have \(\$e\)

OpenStudy (anonymous):

45.45 years